# Constraint-Aware Discrete-Time PID Gain Optimization for Robotic Joint Control Under Actuator Saturation

### v2

@@ Abstract
The precise regulation of rotary actuation is fundamental in autonomous robotics, yet practical PID loops deviate from continuous-time theory due to discrete-time execution, actuator saturation, and small delays and measurement imperfections. We present an implementation-aware analysis and tuning workflow for saturated discrete-time joint control. We (i) derive PI stability regions under Euler and exact zero-order-hold (ZOH) discretizations using the Jury criterion, (ii) evaluate a discrete back-calculation anti-windup realization under saturation-dominant regimes, and (iii) propose a hybrid-certified Bayesian optimization workflow that screens analytically unstable candidates and behaviorally unsafe transients while optimizing a robust IAE objective with soft penalties on overshoot and saturation duty. Baseline sweeps ($\tau=1.0$ s, $\Delta t=0.01$ s, $u\in[-10,10]$) quantify rise/settle trends for P/PI/PID. Under a randomized model family emulating uncertainty, delay, noise, quantization, and tighter saturation, robustness-oriented tuning improves median IAE from $0.843$ to $0.430$ while keeping median overshoot below $2\%$. In simulation-only tuning, the certification screen rejects $11.6\%$ of randomly sampled gains within bounds before full robust evaluation, improving sample efficiency without hardware experiments.
@@ 1 Introduction
The precise regulation of rotary actuation constitutes a fundamental challenge in the engineering of autonomous robotic systems, ranging from industrial manipulators to self-guided electric vehicle charging mechanisms. In these cyber-physical domains, the fidelity of motion control determines the operational success of the system. The Proportional-Integral-Derivative (PID) controller remains the preeminent standard in feedback control systems due to its versatile structure and proven capacity to ensure asymptotic stability.
@@ 2.2 Research objectives
Table 1 lists the key symbols used throughout this paper.
@@ 4 Design and methodology
This section separates the design/methods (models, discretizations, constraints, objectives, and workflows) from the experimental results presented later. This organization follows the mentor feedback to avoid mixing design derivations with result interpretation.
@@ 4.1.5 Actuator saturation, derivative filtering, and discrete anti-windup
The derivative term is well known to amplify measurement noise; practical PID implementations therefore use a filtered derivative (band-limited differentiator) rather than an ideal discrete difference [1, 2]. In our simulation pipeline, the derivative estimate is passed through a first-order low-pass filter; this does not change the conceptual derivations, but it yields behavior closer to embedded practice in the delay/noise experiments.
To explicitly model practical embedded control, we include a discrete back-calculation anti-windup augmentation:
@@ 4.3.2 Exact ZOH discretization
4.4 Saturation-dominant behavior and anti-windup (design)
To evaluate saturation and windup directly (rather than only in the mostly-linear baseline sweeps), we define a saturation-dominant scenario by (i) tightening the actuator clamp so that early transients saturate, and (ii) adding a small input delay and lightly damped second-order actuator dynamics to emulate realistic servo behavior. In this regime, integral windup is expected to elongate recovery tails unless anti-windup is applied. The experiment reports both the output trajectory and the commanded/saturated effort to directly visualize clamping and recovery.
4.5 Robust constraint-aware gain tuning (simulation-only design)
Without access to a physical robot, we emulate realistic non-idealities by evaluating gains across a randomized family of models:
@@ 4.5 Robust constraint-aware gain tuning (simulation-only design)
Parameterization. Unless noted, we set $\%\mathrm{OS}_{\max}=5\%$, $(\lambda_{os},\lambda_{sat},\lambda_{u})=(1,5,0.5)$, and compute all metrics over a 22 s horizon with unit-step reference; $\mathrm{IAE}$ is normalized by the horizon length to keep $J$ dimensionless. where $\mathcal{M}$ indexes randomized models and $\mathrm{sat\_duty}_{m}$ is the fraction of time the actuator saturates. This objective is compatible with Bayesian optimization (BO) because it is expensive, noisy, and non-convex.
4.6 ML-aided gain tuning via Bayesian optimization (workflow)
Manual tuning requires iterative trial-and-error. An alternative is to treat the closed-loop simulation as a black box and optimize gains by minimizing a scalar cost function built from (9)–(13) (or the robust objective in (21)). Bayesian optimization iteratively proposes $(K_{p},K_{i},K_{d})$, evaluates $J$, and updates a surrogate model to select the next candidate [11, 10].
Figure 3: BO workflow: propose gains, simulate, score $J$, update surrogate.
4.6.1 Hybrid certification for safe gain exploration (key novelty)
A central risk of black-box gain tuning is evaluating controllers that are numerically stable under saturation yet dynamically unacceptable (e.g., violent overshoot, sustained saturation, or long windup recovery). We introduce a hybrid certification filter that prevents such evaluations while retaining sample efficiency:
@@ 4.6.1 Hybrid certification for safe gain exploration (key novelty)
  • Behavioral certification (safety/performance): we additionally require that a short-horizon nominal “robotic” check on a second-order actuator model satisfies bounded overshoot and does not remain saturated for the entire transient. This rejects aggressively destabilizing gains before expensive robust evaluation.
This hybrid filter is intentionally lightweight and simulation-only: it does not assume access to hardware, yet it enforces robotics-relevant behavior constraints rather than merely preventing divergence.
@@
4.6.2 Actuator-realistic second-order joint model (robotics benchmark)
The first-order plant in (1) captures a velocity-limited actuator, but many robotic joints exhibit second-order behavior dominated by inertia and damping. To emulate this without hardware, we evaluate gains on the canonical normalized second-order actuator model
@@ 4.6.2 Actuator-realistic second-order joint model (robotics benchmark)
where $\theta$ is joint position, $(\omega_{n},\zeta)$ set natural frequency and damping, $K_{u}$ maps input to torque/acceleration, and $d(t)$ is a disturbance torque. We augment (22) with practical non-idealities: (i) input delay of 1–2 samples, (ii) Coulomb and viscous friction, (iii) a small deadzone, (iv) encoder quantization and additive noise, and (v) tight saturation that makes anti-windup essential. This model bridges the gap between an educational first-order surrogate and actuator behavior common in robotics.
4.6.3 Expanded robust benchmark suite (no hardware required)
@@ 5 Experiments and results
This section reports baseline gain sweeps and then progressively more realistic simulation results (saturation-dominant tests, Monte Carlo robustness, and expanded-benchmark BO results). The design/derivation details appear earlier in Sec. 4.
5.1 Simulation validity and why it is informative for robotics tuning
Robotics control studies increasingly rely on benchmark-grade simulation when hardware access is limited or when exhaustive stress-testing on physical systems is impractical. Modern suites emphasize high-throughput evaluation under uncertainty (rather than a single nominal plant) to quantify robustness and safety-relevant outliers [16, 17, 21]. We adopt the same philosophy: the claims in this paper are comparative—they describe which tuning procedures are safer and more sample-efficient under the stated non-idealities—not that the absolute numerical metrics will match any specific robot without identification.
Our simulation pipeline is designed to be implementation-faithful in the aspects that most strongly affect discrete-time PID behavior:
  • Digital loop semantics: all controllers are executed in discrete time with ZOH actuation and explicit effort clamping, matching embedded joint loops.
  • Non-idealities that dominate PID failure modes: we inject sample delay, measurement noise, and encoder quantization (Sec. 4.5), and we evaluate under tight saturation where windup is frequent (Sec. 4.4).
  • Dynamics richness beyond a single pole: in addition to the interpretable first-order surrogate, we evaluate on a lightly damped second-order actuator family with friction and deadzone (Sec. 4.6.2).
Why this supports deployment decisions. Controller selection in practice is constrained by limited tuning budgets and safety risks during exploration. By evaluating candidate gains across a randomized model family, we directly measure (i) the probability of unsafe behavior, (ii) the tail of the performance distribution, and (iii) how quickly an automated tuner converges to a robust solution. These quantities are actionable even without perfect plant identification because they compare procedures under identical assumptions and evaluation budgets.
Table 2: Modeled non-idealities used to emulate embedded joint control without hardware.
@@ 5.1 Simulation validity and why it is informative for robotics tuning
Threats to validity and mitigations. Simulation cannot capture every unmodeled effect (flexibility, temperature drift, saturation nonlinearities beyond clamping). We mitigate this by (i) reporting uncertainty bounds and non-ideality models explicitly, (ii) emphasizing distributional metrics (quantiles and violation rates) rather than single-run traces, and (iii) evaluating on both an interpretable surrogate and a richer actuator family. When desired, the same Safe-BO loop can be plugged into a rigid-body physics engine (e.g., PyBullet) to test full robot dynamics without changing the tuning methodology [18, 19].
@@ 5.3.3 Full PID control
Figure 9: PID step responses at $K_{p}=3.0$, $K_{i}=1.0$, varying $K_{d}$ (merged to save space).
@@ 5.4 Saturation-dominant results and anti-windup
Figs. 1112 report the saturation-dominant scenario described in Sec. $4.$ This test intentionally forces the actuator to clamp during the early transient so that the integrator would otherwise accumulate error while the input is pinned at its limit.
@@ 5.6 Expanded-benchmark Safe-BO results
We now report results for the hybrid-certified Safe-BO workflow described in Sec. 4.6 when evaluated on the expanded benchmark suite (first-order ZOH family + second-order actuator family). This experiment answers two practical questions: (i) does the certification filter prevent clearly unacceptable evaluations while still allowing progress, and (ii) does the constrained workflow remain sample-efficient compared to unconstrained search?
@@ 6 Discussion
Baseline sweeps confirm that for a delay-free first-order plant, $K_{p}$ primarily sets the steady-state tail (but cannot remove steady-state error), and $K_{i}$ dominates tail removal and IAE reduction. On this nominal single-pole system, derivative action yields only modest shaping; however, once delay and higher-order actuator dynamics are introduced, $K_{d}$ becomes materially beneficial for damping and transient control. The stability-region derivations provide a fast guardrail for PI gains before simulation, and the saturation-dominant tests demonstrate why anti-windup is essential once the clamp is active. Robustness-oriented tuning and Safe-BO results illustrate a practical workflow for simulation-only tuning: optimizing gains against randomized model families improves typical-case tracking while keeping overshoot bounded and rejecting a large fraction of unsafe candidate gains.
@@ 6.1 Limitations and threats to validity
The results emphasize that “implementation details”—sampling, saturation, and small delays—change both the safe gain set and the qualitative closed-loop behavior, even for otherwise simple plants. The ZOH and Euler PI regions provide an interpretable first-line guardrail for the integral action that most strongly governs steady-state and tail behavior.
Scope. The analytic guard is nominal and PI-specific; for full PID tuning we therefore rely on the behavioral screen and robust evaluation, and we report simulation-based evidence rather than hardware guarantees. We explicitly model uncertainty, delay, noise, and quantization to approximate common embedded joint loops, but the benchmark models remain simplified abstractions.
@@ 7 Conclusion
This paper reformulates a classic PID tuning problem in the explicitly practical setting of discrete-time embedded control with actuator limits. In addition to baseline P/PI/PID characterization, we derive PI stability regions using the Jury criterion, demonstrate saturation-dominant anti-windup behavior, and propose a simulation-only robustness-oriented tuning workflow that emulates key hardware non-idealities through a randomized model family. We further introduce a hybrid-certified Safe-BO workflow to prevent unsafe gain exploration and demonstrate its behavior on a second-order actuator benchmark suite.
@@ B.1 Derivative filtering and noise
implemented in discrete time using a stable recursion. In our simulation studies, adding measurement noise shifts the trade-off between overshoot and IAE, and derivative action becomes beneficial only when the effective bandwidth is chosen relative to the sampling period and noise level. As a practical guideline, we recommend selecting $\omega_{f}$ such that $\omega_{f}T_{s}\in[0.1,\,0.3]$ and evaluating robustness under noise/quantization before increasing $K_{d}$.
@@ B.3 Delay, quantization, and discrete-time guards
Computational delay and quantization act like additional dynamics that reduce phase margin. We therefore propose using analytic discrete-time stability guards (Jury conditions) as a prefilter for any automated tuning loop, and then validating candidates under randomized uncertainty (delay, quantization, noise) before declaring a gain set “robust”.
@@ B.5 Objective design for tuning
Robotics PID tuning is frequently framed as “make tracking good,” but the objective choice strongly influences the resulting gains. A pure tracking-error objective tends to push gains toward aggressive behavior that exploits saturation, whereas adding explicit penalties on effort and effort rate discourages chattering and reduces thermal stress. In this work, we evaluate candidates under a composite objective that includes IAE, overshoot, and a saturation-duty term; this combination was empirically the most predictable trade-off: strong penalties on persistent saturation and overshoot tend to produce well-damped transients without inflating control effort.
@@ B.6 Reporting standards for reproducible tuning
To support reproducibility, each reported experiment specifies: sampling period $T_{s}$, saturation bounds, noise/quantization model, simulation horizon, and the exact objective functional. We recommend reporting all five in any robotics PID study to enable apples-to-apples comparisons.
@@ Appendix D Hybrid-certified Bayesian optimization procedure
    Analytic prefilter: reject any candidate that violates the discrete-time stability guards derived for the nominal model (Jury conditions for PI, and the corresponding linearized constraints for PID).
@@
    Behavioral check: run a short-horizon simulation on a small subset of uncertain models. Reject candidates that produce divergence, excessive overshoot, or persistent saturation.

### v3

@@ Abstract
The precise regulation of rotary actuation is fundamental in autonomous robotics, yet practical PID loops deviate from continuous-time theory due to discrete-time execution, actuator saturation, and small delays and measurement imperfections. We present an implementation-aware analysis and tuning workflow for saturated discrete-time joint control. We (i) derive PI stability regions under Euler and exact zero-order-hold (ZOH) discretizations using the Jury criterion, (ii) evaluate a discrete back-calculation anti-windup realization under saturation-dominant regimes, and (iii) propose a hybrid-certified Bayesian optimization workflow that screens analytically unstable candidates and behaviorally unsafe transients while optimizing a robust IAE objective with soft penalties on overshoot and saturation duty. Baseline sweeps ($\tau=1.0$ s, $\Delta t=0.01$ s, $u\in[-10,10]$) quantify rise/settle trends for P/PI/PID. Under a randomized model family emulating uncertainty, delay, noise, quantization, and tighter saturation, robustness-oriented tuning improves median IAE from $0.687$ to $0.470$ while keeping median overshoot below $2\%$. In simulation-only tuning, the certification screen rejects $11.6\%$ of randomly sampled gains within bounds before full robust evaluation, improving sample efficiency without hardware experiments.
@@ 1 Introduction
The precise regulation of rotary actuation constitutes a fundamental challenge in the engineering of autonomous robotic systems, ranging from industrial manipulators to self-guided electric vehicle charging mechanisms. In these cyber-physical domains, the fidelity of motion control determines the operational success of the system. The Proportional-Integral-Derivative (PID) controller remains a widely used standard in feedback control systems due to its versatile structure and proven capacity to ensure asymptotic stability.
@@ 4 Design and methodology
This section specifies the plant and actuator models, discrete-time implementation, saturation handling, and evaluation metrics used throughout the paper. It then derives discrete-time PI stability regions via the Jury criterion under both Euler and exact ZOH discretizations and formalizes the constrained objective used for robust gain selection. These components define the feasibility and scoring framework used in the baseline sweeps, saturation-dominant experiments, Monte Carlo robustness evaluation, and constrained Bayesian optimization results.
@@ 4.1.5 Actuator saturation, derivative filtering, and discrete anti-windup
The derivative term is well known to amplify measurement noise; practical PID implementations therefore use a filtered derivative (band-limited differentiator) rather than an ideal discrete difference [1, 2]. In our pipeline, the derivative estimate is passed through a first-order low-pass filter; this does not change the conceptual derivations, but it yields behavior closer to embedded practice in the delay/noise experiments.
To explicitly model embedded control under saturation, we include a discrete back-calculation anti-windup augmentation:
@@ 4.3.2 Exact ZOH discretization
4.4 Saturation-dominant behavior and anti-windup
To evaluate saturation and windup directly (rather than only in the baseline sweeps), we define a saturation-dominant scenario by (i) tightening the actuator clamp so that early transients saturate, and (ii) adding a small input delay and lightly damped second-order actuator dynamics to emulate realistic servo behavior. In this regime, integral windup is expected to elongate recovery tails unless anti-windup is applied. The experiment reports both the output trajectory and the commanded/saturated effort to directly visualize clamping and recovery.
4.5 Robust constraint-aware gain tuning
To quantify robustness under non-idealities, we evaluate gains across a randomized family of models:
@@ 4.5 Robust constraint-aware gain tuning
Parameterization. Unless noted, we set $\%\mathrm{OS}_{\max}=5\%$, $(\lambda_{os},\lambda_{sat},\lambda_{u})=(1,5,0.5)$, and compute all metrics over a $2$ s horizon with unit-step reference; $\mathrm{IAE}$ is normalized by the horizon length to keep $J$ dimensionless. where $\mathcal{M}$ indexes randomized models and $\mathrm{sat\_duty}_{m}$ is the fraction of time the actuator saturates. This objective is compatible with Bayesian optimization (BO) because it is expensive, noisy, and non-convex.
4.6 ML-aided gain tuning via Bayesian optimization
Manual tuning requires iterative trial-and-error. An alternative is to treat the closed-loop evaluation as a black box and optimize gains by minimizing a scalar cost function built from (9)–(13) (or the robust objective in (21)). Bayesian optimization iteratively proposes $(K_{p},K_{i},K_{d})$, evaluates $J$, and updates a surrogate model to select the next candidate [11, 10].
Figure 3: BO workflow: propose gains, evaluate, score $J$, update surrogate.
4.6.1 Hybrid certification for safe gain exploration
A central risk of black-box gain tuning is evaluating controllers that are numerically stable under saturation yet dynamically unacceptable (e.g., extreme overshoot, sustained saturation, or long windup recovery). We introduce a hybrid certification filter that prevents such evaluations while retaining sample efficiency:
@@ 4.6.1 Hybrid certification for safe gain exploration
  • Behavioral certification (constraints): we additionally require that a short-horizon nominal check on a second-order actuator model satisfies bounded overshoot and does not remain saturated for the entire transient. This rejects aggressively destabilizing gains before expensive robust evaluation.
This filter is lightweight and evaluation-driven: it enforces robotics-relevant behavior constraints rather than merely preventing divergence.
@@
4.6.2 Actuator-realistic second-order joint model
The first-order plant in (1) captures a velocity-limited actuator, but many robotic joints exhibit second-order behavior dominated by inertia and damping. To emulate this, we evaluate gains on the normalized second-order actuator model
@@ 4.6.2 Actuator-realistic second-order joint model
where $\theta$ is joint position, $(\omega_{n},\zeta)$ set natural frequency and damping, $K_{u}$ maps input to torque/acceleration, and $d(t)$ is a disturbance torque. We augment (22) with practical non-idealities: (i) input delay of 1–2 samples, (ii) Coulomb and viscous friction, (iii) a small deadzone, (iv) encoder quantization and additive noise, and (v) tight saturation that makes anti-windup essential. This model bridges a single-pole surrogate and actuator behavior commonly observed in robotics.
4.6.3 Expanded robust benchmark suite
@@ 5 Experiments and results
This section reports baseline gain sweeps and then progressively more realistic evaluation results (saturation-dominant tests, Monte Carlo robustness, and expanded-benchmark BO results).
5.1 Evaluation validity and scope
The results are reported as comparative outcomes under a consistent evaluation protocol. All controllers are executed in discrete time with ZOH actuation and explicit effort clamping, and all candidate gains are tested against non-idealities that commonly dominate PID performance: sample delay, measurement noise, encoder quantization, and tighter saturation limits (Sec. 4.5). In addition to a first-order surrogate, gains are evaluated on a lightly damped second-order actuator family with friction and deadzone (Sec. 4.6.2). These evaluations quantify typical behavior and violation rates under uncertainty rather than relying on a single nominal trajectory.
Table 2: Modeled non-idealities used to emulate embedded joint control.
@@ 5.3.3 Full PID control
Figure 9: PID step responses at $K_{p}=3.0$, $K_{i}=1.0$, varying $K_{d}$ (merged).
@@ 5.4 Saturation-dominant results and anti-windup
Figs. 1112 report the saturation-dominant scenario described in Sec. 4.4. This test forces the actuator to clamp during the early transient so that, without compensation, the integrator would otherwise accumulate error while the input is pinned at its limit.
@@ 5.6 Expanded-benchmark Safe-BO results
We now report results for the hybrid-certified Safe-BO workflow described in Sec. 4.6 when evaluated on the expanded benchmark suite (first-order ZOH family + second-order actuator family). This experiment addresses two practical questions: (i) does the certification filter prevent clearly unacceptable evaluations while still allowing progress, and (ii) does the constrained workflow remain sample-efficient compared to unconstrained search?
@@ 6 Discussion
Baseline sweeps confirm that for a delay-free first-order plant, $K_{p}$ primarily sets the steady-state tail (but cannot remove steady-state error), and $K_{i}$ dominates tail removal and IAE reduction. On this nominal single-pole system, derivative action yields only modest shaping; however, once delay and higher-order actuator dynamics are introduced, $K_{d}$ becomes materially beneficial for damping and transient control. The stability-region derivations provide a fast guardrail for PI gains before evaluation, and the saturation-dominant tests demonstrate why anti-windup is essential once the clamp is active. Robustness-oriented tuning and Safe-BO results illustrate a practical workflow for constrained tuning: optimizing gains against randomized model families improves typical-case tracking while keeping overshoot bounded and rejecting a large fraction of unsafe candidate gains.
@@ 6.1 Limitations and threats to validity
The results emphasize that sampling, saturation, and small delays change both the safe gain set and the qualitative closed-loop behavior, even for simple plants. The ZOH and Euler PI regions provide an interpretable first-line guardrail for the integral action that most strongly governs steady-state and tail behavior.
Scope. The analytic guard is nominal and PI-specific; for full PID tuning we therefore rely on the behavioral screen and robust evaluation, and we report evaluation-based evidence rather than hardware guarantees. We explicitly model uncertainty, delay, noise, and quantization to approximate common embedded joint loops, but the benchmark models remain simplified abstractions.
@@ 7 Conclusion
This paper reformulates a classic PID tuning problem in the practical setting of discrete-time embedded control with actuator limits. In addition to baseline P/PI/PID characterization, we derive PI stability regions using the Jury criterion, demonstrate saturation-dominant anti-windup behavior, and propose a robustness-oriented tuning workflow that emulates key non-idealities through a randomized model family. We further introduce a hybrid-certified Safe-BO workflow to prevent unsafe gain exploration and demonstrate its behavior on a second-order actuator benchmark suite.
@@ B.1 Derivative filtering and noise
implemented in discrete time using a stable recursion. In our evaluation studies, adding measurement noise shifts the trade-off between overshoot and IAE, and derivative action becomes beneficial only when the effective bandwidth is chosen relative to the sampling period and noise level. As a practical guideline, we recommend selecting $\omega_{f}$ such that $\omega_{f}T_{s}\in[0.1,\,0.3]$ and evaluating robustness under noise/quantization before increasing $K_{d}$.
@@ B.3 Delay, quantization, and discrete-time guards
Computational delay and quantization act like additional dynamics that reduce phase margin. We therefore propose using analytic discrete-time stability guards (Jury conditions) as a prefilter for any automated tuning loop, and then validating candidates under randomized uncertainty (delay, quantization, noise) before declaring a gain set robust.
@@ B.5 Objective design for tuning
Robotics PID tuning is frequently framed as “make tracking good,” but the objective choice strongly influences the resulting gains. A pure tracking-error objective tends to push gains toward aggressive behavior that exploits saturation, whereas adding explicit penalties on effort and effort rate discourages chattering and reduces thermal stress. In this work, we evaluate candidates under a composite objective that includes IAE, overshoot, and a saturation-duty term; this combination yields a predictable trade-off between tracking performance and saturation exposure.
@@ Appendix D Hybrid-certified Bayesian optimization procedure
    Analytic prefilter: reject any candidate that violates the discrete-time stability guards derived for the nominal model (Jury conditions for PI, and corresponding linearized constraints for PID).
@@
    Behavioral check: run a short-horizon evaluation on a small subset of uncertain models. Reject candidates that produce divergence, excessive overshoot, or persistent saturation.